# Improving LLM-based Machine Translation with Systematic Self-Correction

### v2

@@ Abstract
Large Language Models (LLMs) have achieved impressive results in Machine Translation (MT). However, careful evaluations by human reveal that the translations produced by LLMs still contain multiple errors. Importantly, feeding back such error information into the LLMs can lead to self-correction and result in improved translation performance. Motivated by these insights, we introduce a systematic LLM-based self-correcting translation framework, named TER, which stands for Translate, Estimate, and Refine, marking a significant step forward in this direction. Our findings demonstrate that 1) our self-correction framework successfully assists LLMs in improving their translation quality across a wide range of languages, whether it’s from high-resource languages to low-resource ones or whether it’s English-centric or centered around other languages; 2) TER exhibits superior systematicity and interpretability compared to previous methods; 3) different estimation strategies yield varied impacts on AI feedback, directly affecting the effectiveness of the final corrections. We further compare different LLMs and conduct various experiments involving self-correction and cross-model correction to investigate the potential relationship between the translation and evaluation capabilities of LLMs. [^1] We release our data and code at https://github.com/fzp0424/self_correct_mt
Improving LLM-based Machine Translation with Systematic Self-Correction
@@ 1 Introduction
The results of the General Machine Translation Task  [^2] WMT 2023 Shared Task: General Machine Translation in WMT23 indicate that LLM-based machine translation systems  (Agrawal et al., 2023; Vilar et al., 2023; Zhang et al., 2023; He et al., 2023; Gao et al., 2023; Garcia et al., 2023; Wu and Hu, 2023; Hendy et al., 2023; Peng et al., 2023a; Moslem et al., 2023; Zhu et al., 2023; Liang et al., 2023), especially GPT-4  (Achiam et al., 2023) using few-shot prompting, have taken top positions in the majority of subtasks (Kocmi et al., 2023). However, taking the whole 1976 test pairs from the WMT23 Zh-En dataset as a probing study, even with GPT-4 (the best submission in WMT23), only 332 pairs achieved a perfect score (i.e., no errors were identified upon manual inspection) according to the Multi-dimensional Quality Metrics (MQM, Freitag et al., 2021)   [^3] https://github.com/google/wmt-mqm-human-evaluation . Interestingly, when we feed back the information from the human MQM evaluations into GPT-4, asking it to correct its initial translation based on this feedback, we observe that the errors can be corrected and there is also an improvement in the metric score, as shown in Figure 1. Such findings inspire us to leverage evaluative feedback on translations as a means to facilitate the correction and improvement of initial translations by LLMs.
@@
Involving LLMs to automatically do the self refinement is increasingly gaining attention (Saunders et al., 2022; Welleck et al., 2022; Pan et al., 2023; Madaan et al., 2023; Shinn et al., 2023; Chen et al., 2023b; Li et al., 2023; Gou et al., 2023). In the field of Machine Translation (MT), Chen et al. (2023a) investigate the use of LLMs to rewrite translations by feeding back the previous translations, achieving changes at the lexical and structural levels while maintaining translation quality. Raunak et al. (2023) ask GPT-4 to refine translations with the inclusion of the edit proposals produced by Chain-of-Thought (CoT) (Kojima et al., 2022) strategy, indicating that GPT-4 is adept at translation post-editing. However, these efforts encounter several challenges: 1) There is a lack of clear assessment of the original translations, with evaluations either absent or merged into the correction process; 2) The feedback provided is ineffective as it lacks clear guidance or contains redundant information, leading to difficulties in comprehension by LLMs; 3) The initial translations originate from external sources, yet the capability of LLMs for self-correction remains uninvestigated; 4) The translation quality improvements are not strong, and in some scenarios, a decline has been noted.
@@
In this paper, we propose an LLM-based self-correcting translation framework, termed TER: Translate, Estimate, and Refine. Translate module utilizes an LLM for translation, ensuring that our translations are internally sourced. Estimate module receives the initial translation and provides systematic estimations of the translation quality. This assessment serves as feedback. Refine module performs translation corrections based on the information from the preceding two modules. Our experimental results indicate that our approach is more effective in improving translation quality compared to the baselines, across both high- and low-resource languages, as well as English-centric and non-English-centric direct translations, especially in semantic-related metrics (COMET (Rei et al., 2020), COMETKiwi (Rei et al., 2022), and BLEURT (Sellam et al., 2020)). We also explore the combined performance of our three modules under various prompting strategies in self-correction and cross-model correction. Additionally, we conduct a more detailed analysis of the estimation component according to types of translation errors. Our contributions are as follows:
  • We propose the first LLM-based self-correcting translation framework, TER, featuring a standalone estimation module and establishing a feedback concept within self-correcting MT, weaving into an interpretable translation system.
  • TER significantly improves translation quality across various language pairs with different LLMs, e.g., ChatGPT (Achiam et al., 2023), Gemini (Team et al., 2023), Claude (Anthropic, 2023), demonstrating effectiveness for both high-resource and low-resource languages as well as English-centric and non-English-centric translations.
  • We conduct a comprehensive analysis to provide clear insights into how different components of TER interact and contribute to overall performance, facilitating targeted improvements and innovations in the field. We demonstrate that translation ability and quality estimation ability exhibit similar trends in specific language pairs, while error correction ability complements them in other language pairs, providing novel insights for future work.
@@
Figure 2: TER framework includes three steps: Translate, Estimate, and Refine. All steps are executed using different prompts ($\mathcal{T}_{translate}$, $\mathcal{T}_{estimate}$, $\mathcal{T}_{refine}$). We detail our prompting strategies in Section 3.
Figure 3: Results for 16 translation directions using GPT-3.5-turbo. IT: the initial translation using few-shot prompt; SCoT: Structured Chain-of-Thought (Raunak et al., 2023); CT: inserting the word ${}^{\prime\prime}$${}^{\prime\prime}$ to do the contrastive translation (Chen et al., 2023a).
Figure 4: Results for human preference study, comparing TER with IT, SCoT, and CT. The data for the first row of subfigures comes from WMT22 tested on GPT-3.5-turbo, while the experiments for the second row of subfigures were conducted on our WMT23 Zh-En dataset using three models (GPT-3.5-turbo, Gemini-Pro, Claude-2).
2 TER
@@
Figure 2 illustrates the workflow of TER. The TER framework consists of three modules: Translate, Estimate, and Refine. Given an LLM $\mathcal{M}$ with fixed parameters $\theta$ and a language pair $\mathcal{X}$-$\mathcal{Y}$ (source-target), we first utilize $\mathcal{M}$ to generate the initial translation for segment $x$ with the prompt $\mathcal{T}_{translate}$. Got the initial translation $y$, we use the same LLM $\mathcal{M}$ with prompt $\mathcal{T}_{estimate}$ for estimation. Using the estimation as feedback, we leverage $\mathcal{M}$ to refine the initial translation with $\mathcal{T}_{refine}$.
@@
Quality Estimation (QE) in the field of MT refers to a method predicting the quality of a given translation rather than assessing how similar it is to a reference segment. Nonetheless, the learned models (Rei et al., 2022; Perrella et al., 2022; Gowda et al., 2023; Juraska et al., 2023; Guerreiro et al., 2023) focus on offering a single numerical score to gauge the quality. Even when they can predict spans of errors, they only provide information on the level of severity. This presents three main challenges: 1) The first is incorporating additional models causes inconvenience and requires a significant amount of annotated data for training; 2) The second is understanding the meaning behind the numerical representation; 3) The last is obtaining the feedback with guiding significance. Gold MQM annotation requires expensive human efforts, which runs counter to automatic QE. Recently, Kocmi and Federmann (2023) and Fernandes et al. (2023) successfully leverage LLMs to automatically annotate with MQM guideline (Freitag et al., 2021). Guerreiro et al. (2023) also showcase that using LLMs to assess translation quality can achieve performance comparable to models trained on extensive annotated datasets. Prompting with $\mathcal{T}_{estimate}$, we can make a black-and-white judgment on whether refinement is needed, rather than empirically using the learned metrics and setting a numerical threshold. Besides, when doing refinement, the result generated during the estimating process also provides important guidance.
@@
Large language models have elevated the level of machine translation to unprecedented heights. It is relatively easy to progress from a passing level to excellence, whereas reaching the pinnacle from excellence is no easy feat. Chen et al. (2023a) and Raunak et al. (2023) have shown the potential of leveraging LLMs to refine the translation. However, both of them lack an explicit estimation module and feedback concept, which would make this improvement process unclear. Therefore, we collect the estimation as the feedback to establish $\mathcal{T}_{refine}$ to refine the initial translation.
In a word, our TER self-correction framework requires only one LLM to accomplish all the functionalities of the aforementioned modules, providing insights into the self-evolution of LLM-based translation systems.
@@
Table 1: Comparison of quality improvement for different variants of the TER modules in our sampled WMT23 Zh-En dataset. $\mathcal{T}_{refine}-\alpha$ uses only the feedback from the Estimate module, $\mathcal{T}_{refine}-\beta$ uses few-shot exemplars in the Translate module and the feedback. Bold results indicate the best in each section.
@@ 3 Experimental Setup
In this paper, we accessed all LLMs through their APIs including GPT-3.5-turbo  [^5] We utilize gpt-3.5-turbo-0613 whose training data is up to date as of September 2021 https://platform.openai.com/docs/models/gpt-3-5, Claude-2  [^6] https://www.anthropic.com/index/claude-2 and Gemini-Pro  [^7] https://cloud.google.com/vertex-ai/docs/generative-ai/learn/models. Due to the limited computational resources, we did not include open-source LLMs.
Figure 5: COMET scores for involving various feedback estimation strategies in the TER. "-" denotes the baseline where no estimation is involved, representing the initial translation (IT). zero-shot and few-shot reflect the use of different prompting methods with GPT-3.5-turbo, while GPT-4 w/ human indicates estimations made using GPT-4 with human assistance.
@@
Table 2: Pairwise human evaluation on zero-shot and few-shot estimation.
Prompting Strategies
For $\mathcal{T}_{translate}$, we follow the setting in Hendy et al. (2023) and Kocmi et al. (2023), using the few-shot (5-shot) high-quality translation pairs. For $\mathcal{T}_{estimate}$, we use few-shot (3-shot) multi-lingual MQM annotated examples to ask LLMs to estimate the initial target in a reference-free scenario. Kocmi and Federmann (2023) have demonstrated that this prompt method ensures the estimation can be executed across any language pairs and can rival metrics models trained on a large amount of MQM annotated data. As for $\mathcal{T}_{refine}$, $\mathcal{T}_{translate}-\alpha$ represents only using the feedback, while $\mathcal{T}_{translate}-\beta$ adheres estimation feedback to the few-shot translation examples. All prompts are detailed in Appendix G.
@@
To evaluate the quality of a translation, we utilize commonly used metrics and human preference tests.
@@ 4 Results
We compare TER with three baselines, i.e., 1) the initial translation with few-shot prompting (IT); 2) post-editing with Structured-CoT (SCoT), which uses the MQM annotation instructions (Freitag et al., 2021) to produce the intermediate CoT in the form of an MQM annotation over the source-translation pair (Raunak et al., 2023); 3) Contrastive Translate (CT), which inserts the word “bad” to hint that the previous translation is of low quality, regardless of its actual quality, to form a contrastive prompt (Chen et al., 2023a).
Figure 3 displays our test dataset results across 9 languages and 16 translation directions using GPT-3.5-turbo. Compared to self-translated IT, TER significantly outperforms them in all directions, with an average improvement of 2.48. Notably, the Ru-En pair sees a significant leap of 5.14, and the En-De pair has an impressive boost of 6.88. TER consistently outperforms several post-editing methods across all directions. In terms of BLEU scores, TER leads SCoT by 4.75 and CT by 1.0 on average. For the COMET metric, TER’s performance is ahead of SCoT by 2.24 and beats CT by 0.71. With COMETKiwi, TER surpasses SCoT by 2.10 and CT by 0.68. On the Bleurt metric, it exceeds SCoT by 2.48 and CT by 1.11. To confirm the versatility of the TER framework with different models, we evaluated Gemini-Pro and Claude-2 on our WMT23 Zh-En dataset, comparing them against all baselines. TER demonstrated impressive performance in these tests as well (see Table 8).
To further compare the translation quality of our TER method, we conduct pairwise human evaluations in Ja-En, En-Zh, and Zh-En translation directions. Figure 4 demonstrates that when compared with IT, the proportion of evaluators who preferred TER translations were 46.2%, 32.2%, and 42.5% respectively, all significantly exceeding the proportion favoring IT. This illustrates the foundational strength of TER in self-correcting the deficiencies of IT. When compared with SCoT and CT, TER was also generally more preferred by human evaluators. These advantages are also clear with Genmini-Pro and Claude-2.
@@
Table 6: Results of the error-correction capabilities of various LLMs. We conducted cross-correction experiments on our sampled WMT23 datasets by swapping out different models in the Refine module. "-" denotes the baseline without refinement. Bold results indicate the best in each section.

### v3

# TEaR: Improving LLM-based Machine Translation with Systematic Self-Refinement

@@ Abstract
Large Language Models (LLMs) have achieved impressive results in Machine Translation (MT). However, careful evaluations by human reveal that the translations produced by LLMs still contain multiple errors. Importantly, feeding back such error information into the LLMs can lead to self-refinement and result in improved translation performance. Motivated by these insights, we introduce a systematic LLM-based self-refinement translation framework, named TEaR, which stands for Translate, Estimate, and Refine, marking a significant step forward in this direction. Our findings demonstrate that 1) our self-refinement framework successfully assists LLMs in improving their translation quality across a wide range of languages, whether it’s from high-resource languages to low-resource ones or whether it’s English-centric or centered around other languages; 2) TEaR exhibits superior systematicity and interpretability; 3) different estimation strategies yield varied impacts, directly affecting the effectiveness of the final corrections. Additionally, traditional neural translation models and evaluation models operate separately, often focusing on singular tasks due to their limited capabilities, while general-purpose LLMs possess the capability to undertake both tasks simultaneously. We further conduct cross-model correction experiments to investigate the potential relationship between the translation and evaluation capabilities of general-purpose LLMs. Our code and data are available at https://github.com/fzp0424/self_correct_mt.
TEaR: Improving LLM-based Machine Translation with Systematic Self-Refinement
@@ 1 Introduction
The results of the General Machine Translation Task  [^1] WMT 2023 Shared Task: General Machine Translation in WMT23 indicate that LLM-based machine translation systems  (Vilar et al., 2023; He et al., 2023b; Gao et al., 2023; Wu and Hu, 2023; Hendy et al., 2023; Peng et al., 2023a; Moslem et al., 2023; Liang et al., 2023), especially GPT-4  (Achiam et al., 2023) using few-shot prompting, have taken top positions in the majority of subtasks (Kocmi et al., 2023). However, taking the whole 1976 test pairs from the WMT23 Zh-En dataset as a probing study, even with GPT-4 (the best submission in WMT23), only 332 pairs achieved a perfect score (i.e., no errors were identified upon manual inspection) according to the Multi-dimensional Quality Metrics (MQM, Freitag et al., 2021)   [^2] https://github.com/google/wmt-mqm-human-evaluation . Interestingly, when we feed back the information from the human MQM evaluations into GPT-4, asking it to correct its initial translation based on this feedback, we observe that the errors can be corrected and there is also an improvement in the metric score, as shown in Figure 1. These findings inspire us to use evaluation feedback on translations to facilitate the refinement of initial translations within a single LLM.
@@
Involving LLMs to automatically do the self refinement is increasingly gaining attention (Saunders et al., 2022; Pan et al., 2023; Madaan et al., 2023; Shinn et al., 2023; Li et al., 2023). In the field of MT, Chen et al. (2023) investigate the use of LLMs to rewrite translations by feeding back the previous translations, achieving changes at the lexical and structural levels while maintaining translation quality. Raunak et al. (2023) ask GPT-4 to refine translations from other neural machine translation models based on suggestions from the Chain-of-Thought (CoT) strategy (Kojima et al., 2022), indicating that GPT-4 is adept at translation post-editing. However, these efforts encounter several challenges: 1) There is no clear assessment of the original translations; 2) Ineffective feedback, often lacking clear guidance or containing redundant information, making it difficult for LLMs to comprehend; 3) The self-correction capability of LLMs remains uninvestigated, with insignificant translation quality improvements and occasional declines noted in some scenarios.
@@
In this paper, we propose a systematic and effective LLM-based self-refinement translation framework, termed TEaR: Translate, Estimate, and Refine. Translate module utilizes an LLM for translation, ensuring that our translations are internally sourced. Estimate module receives the initial translation and provides systematic estimations of the translation quality. Refine module performs translation refinements based on the information from the preceding two modules. Our experimental results indicate that TEaR is more effective in improving translation quality compared to previous methods, especially in semantic-related metrics like COMET (Rei et al., 2020). We also explore the combined performance of our three modules under various prompting strategies and conduct a more detailed analysis of the estimation component according to types of translation errors. Traditionally, translation and evaluation models operated separately, focusing on singular tasks due to limited capabilities and training data. For example, models like NLLB (Costa-jussà et al., 2022) excelled at translation but couldn’t evaluate translations, while models like COMET (Rei et al., 2020) could evaluate but not translate. However, general-purpose LLMs can perform both tasks simultaneously and follow instructions well, allowing them to refine translations too. We are curious if the translation capability of LLMs (e.g., GPT-3.5-turbo’s superior Zh-En translation compared to Gemini-Pro) correlates with their translation evaluation capability. Our contributions are as follows:
  • We propose TEaR, the first systematic and effective LLM-based self-refinement translation framework, which includes a standalone estimation module and establishes a feedback mechanism within self-refinement MT, creating an interpretable translation system. Our experimental results indicate that deploying TEaR with different LLMs significantly improves translation quality across various language pairs, demonstrating effectiveness for both high-resource and low-resource languages, as well as for English-centric and non-English-centric translations.
  • We conduct a comprehensive analysis to provide clear insights into how different components of the TEaR framework interact and contribute to overall performance, facilitating targeted improvements and innovations in the field. Our detailed statistical analysis of error information before and after refinement indicates that the estimation module is a critical factor influencing the effectiveness of TEaR.
  • Our cross-model correction experiments show that the translation and estimation abilities of general-purpose LLMs exhibit similar trends in specific language pairs, while their error correction abilities complement these trends in other language pairs. This suggests potential correlations between translation-related tasks within these non-translation-specific LLMs, offering insights for future research.
@@
Figure 2: TEaR framework includes three steps: Translate, Estimate, and Refine. All steps are executed using different prompts ($\mathcal{T}_{translate}$, $\mathcal{T}_{estimate}$, $\mathcal{T}_{refine}$). We detail our prompting strategies in Section 3.
2 TEaR
@@
Figure 2 illustrates the workflow of TEaR, consisting of three modules: Translate, Estimate, and Refine. Given an LLM $\mathcal{M}$ and a language pair $\mathcal{X}$-$\mathcal{Y}$ (source-target), we first utilize $\mathcal{M}$ to generate the initial translation for segment $x$ with prompt $\mathcal{T}_{translate}$. Got the initial translation $y$, we use the same LLM $\mathcal{M}$ with $\mathcal{T}_{estimate}$ for estimation. Using the estimation as feedback, we leverage $\mathcal{M}$ to refine the initial translation with $\mathcal{T}_{refine}$.
@@
Quality Estimation (QE) in MT refers to a method predicting the quality of a given translation, typically using Multidimensional Quality Metrics (MQM) datasets (Freitag et al., 2021), where human experts annotate error spans and assign quality scores (see Appendix C). Nonetheless, the learned neural QE models (Rei et al., 2022; Perrella et al., 2022; Gowda et al., 2023; Juraska et al., 2023; Guerreiro et al., 2023) focus on offering a single numerical score to gauge the quality. Even when they can predict spans of errors, they only provide information on the level of severity. This presents three main challenges: 1) Incorporating additional models leads to extra deployment costs and requires a significant amount of annotated data for training; 2) Understanding the meaning behind the numerical representation; 3) Obtaining feedback that provides meaningful guidance.
Several works have demonstrated that powerful LLMs can perform data annotation tasks similarly to human annotators (He et al., 2023a; Kocmi and Federmann, 2023; Fernandes et al., 2023). Inspired by this, we designed a simplified template $\mathcal{T}_{estimate}$ based on MQM annotator guidelines (Freitag et al., 2021) to leverage LLMs for automatically evaluating translation quality in a human-like manner. Prompting with $\mathcal{T}_{estimate}$, LLM can make a black-and-white judgment on whether the correction is needed, rather than using the learned metrics and setting a numerical threshold. Besides, when doing correction, the result generated during the estimating process also provides important guidance.
@@
LLMs have elevated the level of machine translation to unprecedented heights. Chen et al. (2023) and Raunak et al. (2023) have shown the potential of leveraging LLMs to refine the translation. However, both of them lack an explicit estimation module and feedback concept, which would make this improvement process unclear. Therefore, we collect the estimation as the feedback to establish $\mathcal{T}_{refine}$ to refine the initial translation.
Figure 3: Results for 16 translation directions using GPT-3.5-turbo. IT: the initial translation using few-shot prompt; SCoT: Structured Chain-of-Thought (Raunak et al., 2023); CT: inserting the word ′′bad′′ to do the contrastive translation (Chen et al., 2023). TowerInstruct-13B: the state-of-the-art open-source APE model has already been trained on WMT21 and WMT22 data (Alves et al., 2024).
Figure 4: Results for human preference study, comparing TEaR with IT, SCoT, and CT. The data for the first row of subfigures comes from WMT22 tested on GPT-3.5-turbo, while the experiments for the second row of subfigures were conducted on our WMT23 Zh-En dataset using three models (GPT-3.5-turbo, Gemini-Pro, Claude-2).
@@
Table 1: Comparison of quality improvement for different variants of the TEaR modules in our sampled WMT23 Zh-En dataset. $\mathcal{T}_{refine}-\alpha$ uses only the feedback from the Estimate module, $\mathcal{T}_{refine}-\beta$ uses few-shot exemplars in the Translate module and the feedback. Bold results indicate the best in each section.
@@ 3 Experimental Setup
Due to the multilingual support and capabilities of the models, we focused on the effectiveness of TEaR in general-purpose LLMs through their APIs, including GPT-3.5-turbo [^4] We utilize gpt-3.5-turbo-0613 https://platform.openai.com/docs/models/gpt-3-5, Claude-2555https://www.anthropic.com/index/claude-2, and Gemini-Pro [^6] https://cloud.google.com/vertex-ai/docs/generative-ai/learn/models. We also explored the potential of applying TEaR in stronger LLMs and open-source models, as detailed in Appendix E.
Prompting Strategies
To explore the internal performance and interactions of the TEaR framework components more thoroughly, we consider zero-shot and few-shot prompting strategies for $\mathcal{T}_{translate}$ and $\mathcal{T}_{estimate}$. $\mathcal{T}_{refine}-\alpha$ represents using only the estimation feedback, while $\mathcal{T}_{refine}-\beta$ incorporates estimation feedback into the few-shot translation examples. Prompts we used can be found in Appendix I.
Figure 5: COMET scores for involving various feedback estimation strategies in the TEaR. "-" denotes the initial translation (IT). zero-shot and few-shot reflect the use of different prompting methods with GPT-3.5-turbo, while GPT-4 w/ human indicates estimations made using GPT-4 with human assistance.
@@
Table 2: comparison between zero-shot and few-shot estimation methods evaluated by human experts.
@@
To evaluate the quality of a translation, we utilize commonly used metrics (COMET-22 (Rei et al., 2020), COMETKiwi (Rei et al., 2022), BLEURT-20 (Sellam et al., 2020), and SacreBLEU (Post, 2018)) and human preference tests. Details can be found in Appendix B.
@@ 4 Results
We compare TEaR with four baselines, i.e., 1) the initial translation with few-shot prompting (IT); 2) post-editing with Structured-CoT (SCoT), which guides the LLM to suggest improvements based on the original translation (Raunak et al., 2023); 3) Contrastive Translate (CT), which inserts the word “bad” to hint that the previous translation is of low quality, regardless of its actual quality, to form a contrastive prompt (Chen et al., 2023); 4) TowerInstruct-13B, which is SoTA automatic post-editing (APE) open source model trained on 673k translation examples (Alves et al., 2024).
Figure 3 displays our test dataset results across 9 languages and 16 translation directions using GPT-3.5-turbo. Compared to self-translated IT, TEaR significantly outperforms them in all directions, with an average improvement of 2.48. Notably, the Ru-En pair sees a significant leap of 5.14, and the En-De pair has an impressive boost of 6.88. TEaR consistently outperforms several post-editing methods across all directions. In terms of BLEU scores, TEaR leads SCoT by 4.75 and CT by 1.0 on average. For the COMET metric, TEaR’s performance is ahead of SCoT by 2.24 and beats CT by 0.71. With COMETKiwi, TEaR surpasses SCoT by 2.10 and CT by 0.68. On the Bleurt metric, it exceeds SCoT by 2.48 and CT by 1.11. To confirm the versatility of the TEaR framework with different models, we evaluated Gemini-Pro and Claude-2 on our WMT23 Zh-En dataset, comparing them against all baselines. TEaR demonstrated impressive performance in these tests as well (see Table 8).
To further compare the translation quality of our TEaR method, we conduct pairwise human evaluations in Ja-En, En-Zh, and Zh-En translation directions. Figure 4 demonstrates that when compared with IT, the proportion of evaluators who preferred TEaR translations were 46.2%, 32.2%, and 42.5% respectively, all significantly exceeding the proportion favoring IT. This illustrates the foundational strength of TEaR in self-correcting the deficiencies of IT. When compared with SCoT and CT, TEaR was also generally more preferred by human evaluators. These advantages are also clear with Genmini-Pro and Claude-2.
@@
Table 6: Results of the error-correction capabilities of various LLMs. We conducted cross-correction experiments on our sampled WMT23 datasets by swapping out different models in the Refine module. "-" denotes the baseline without refinement. Bold results indicate the best in each section. We report COMET scores of refined cases.